What impression does the iron cube with ribs a = 4cm make on the table? P =?


1
Expert's answer
2021-01-13T11:35:31-0500

By the definition of the pressure, we have:


"P=\dfrac{F}{A}=\dfrac{mg}{a^2},""P=\dfrac{\rho Vg}{a^2}=\dfrac{\rho a^3g}{a^2}=\rho ag,""P=7860\ \dfrac{kg}{m^3}\cdot 0.04\ m\cdot 9.8\ \dfrac{m}{s^2}=3081.12\ Pa."

Answer:

"P=3081.12\ Pa."